OpenStudy (anonymous):

Solve each trig question below for 0

OpenStudy (anonymous):

\[\sqrt{2}\cos x \tan x-\tan x=0\]

OpenStudy (tkhunny):

Factor, factor, factor!

OpenStudy (anonymous):

Okay...how lol, I'm sorry I'm not seeing anything that i can work with

OpenStudy (tkhunny):

Seriously? \(\tan(x)\left(\sqrt{2}\cos(x) - 1\right) = 0\)

OpenStudy (anonymous):

Dude I know I'm retarded with this stuff, thank you for your help. It is much appreciated.

OpenStudy (tkhunny):

What did you get for solutions?

OpenStudy (anonymous):

0 and pi/4

OpenStudy (tkhunny):

Well, 0 is not in the Domain as you have written it. Did you intend to include it?

OpenStudy (anonymous):

Yes, I mistyped the domain.

OpenStudy (tkhunny):

Then, good call!